OpenStudy (anonymous):

can someone help me with simplifying radicals? I don't understand it from my book and need help with these questions. It would help if someone explained. √14q x 2√4q

OpenStudy (anonymous):

√63x^15y^9/7xy^11 can someone help me do this cause i kinda get this one but my answers don't add up right.

OpenStudy (jhannybean):

\[\sqrt{14} * 2\sqrt{4}\] in order to simplify this, you need to see how you can simplify sqrt 14 and sq rt 4 to the maximum

OpenStudy (anonymous):

2√56q i think thats what you want right?

OpenStudy (jhannybean):

yes :)

OpenStudy (jhannybean):

Are you confused on simplifying radical 56?

OpenStudy (anonymous):

yes !

OpenStudy (jhannybean):

okay, so you have have to break the 56 into simpler multipliers so that you can factor them out, can you name some multipliers of 56?

OpenStudy (anonymous):

14 and 4

OpenStudy (jhannybean):

okay, so by rewriting 56 as 14*4, we get \[2\sqrt{14*4q}\] we can further simplify this as \[2\sqrt{7*2*2*2}\]

OpenStudy (jhannybean):

Now for every pair of 2 OFTHE SAME numbers, you can pull one out and multiply it to the 2 outside the radical :)

OpenStudy (anonymous):

So 4√14 ?

OpenStudy (jhannybean):

Mmhmm :)

OpenStudy (anonymous):

what about the q

OpenStudy (jhannybean):

oh, and variables such as q work the same way.So because we don't have a pair of q's, the q will stay inside the radical! :)

OpenStudy (jhannybean):

and WAH LA!~ you have your answer ^_^

OpenStudy (anonymous):

wait so 4q√14 ? or other way?

OpenStudy (anonymous):

never mind i got it

OpenStudy (jhannybean):

oh okay

OpenStudy (jhannybean):

And @jim_thompson5910 can helpyou with the second question you posted :)

jimthompson5910 (jim_thompson5910):

still need help? your question looks answered, but I'm not sure